# Simple fits for the neutrino luminosities from protoneutron star cooling

### v2

@@ VII Counting rate in future detectors
In HK, the main detection process is inverse beta decay (IBD), $\bar{\nu}_{e}\,p\to n\,e^{+}$, where the final-state positron shows up by its Cherenkov radiation. Due to the similarity of the electron and non-electron antineutrino luminosities and spectra in our models during PNS Kelvin-Helmholtz cooling at times $t\gtrsim 1$ s after bounce, flavor conversions are not a major effect and can be neglected in our simplified analysis. Therefore the expected rate can be written as
@@
where $N_{p}=4.96\times 10^{34}$ is the number of protons for a 0.74 Mton Cherenkov detector. Here, we follow Ref. [30] for the limits of integration in $dE_{\nu}$ and we integrate the positron energies above the energy threshold $E_{\rm th,HK}=5~{}\operatorname{MeV}$. We mention that HK is essentially background free.
Refer to captionRefer to caption
Figure 15: Left panel: Simulation results for the average energy $\langle E_{\bar{\nu}_{e}}\rangle$ (blue), rms energy $\sqrt{\langle E_{\bar{\nu}_{e}}^{2}\rangle}$ (orange) and spectral shape parameter $\beta_{\bar{\nu}_{e}}$ (green) from model 1.62-DD2. Right panel: $\bar{\nu}_{e}$-induced counting rate $tR_{{\bar{\nu}}_{e}}$ in HK computed from simulation data (blue), the fit of the counting rate (orange), and the fit of $tL_{{\bar{\nu}}_{e}}$ (dashed line).
To obtain a numerical estimate, we use data from model 1.62-DD2 as an example, but the same analysis is valid for all of the models. In the left panel of Fig. 15 we plot the average energy $\langle E_{\bar{\nu}_{e}}\rangle$, the root mean square (rms) energy $\sqrt{\langle E_{\bar{\nu}_{e}}^{2}\rangle}$, and the shape parameter $\beta_{\bar{\nu}_{e}}$ as a function of the time, between 1 s and $t_{\bar{\nu}_{e},{\rm c}}=7.94\,\operatorname{s}$. These quantities exhibit a weak time dependence, with $\langle E_{\bar{\nu}_{e}}\rangle\approx(12-13.5)\,\operatorname{MeV}$, $\sqrt{\langle E_{\bar{\nu}_{e}}^{2}\rangle}\approx(14-15)\,\operatorname{MeV}$ and $\beta_{\bar{\nu}_{e}}\approx 2.5-3$ at $t\lesssim 5\,\operatorname{s}$ and a decrease at later times.
The rate in HK can be simply estimated as
$R_{\bar{\nu}_{e}}\approx\frac{L_{\bar{\nu}_{e}}}{4\pi\,D^{2}\langle E_{\bar{\nu}_{e}}\rangle}\,N_{\rm HK}\,\langle\sigma\rangle\,,$ (9)
@@
dependent on $\langle E_{\bar{\nu}_{e}}\rangle$ and $\beta_{\bar{\nu}_{e}}$ [and also on $\sqrt{\langle E_{\bar{\nu}_{e}}^{2}\rangle}$ via Eq. (7)]. Therefore, we expect that the observed rate $R_{\bar{\nu}_{e}}$ will follow the time dependence of the neutrino luminosity because of the weak time dependence of the other parameters, and that we can also fit it with our analytical formula in Eq. (2). In the right panel of Fig. 15 we show the quantity $tR_{\bar{\nu}_{e}}$ obtained from Eq. (8) using data from model 1.62-DD2 (blue line), its fit (orange), and the fitted $tL_{{\bar{\nu}}_{e}}$ (dashed line). As shown by the best-fit parameters for $t\,R_{\bar{\nu}_{e}}$ given in the plot, the fit leads to $C=2.72\times 10^{4}\,\operatorname{s}^{-1}$, $\alpha=0.54$, $\tau=6.17\,\operatorname{s}$ and $n=4.61$, to be compared with the ones for $tL_{\bar{\nu}_{e}}$. The parameters $\alpha_{\bar{\nu}_{e}}$, $\tau_{\bar{\nu}_{e}}$ and $n_{\bar{\nu}_{e}}$ are well reconstructed, while one could get information on the normalization factor $C_{\bar{\nu}_{e}}$ by inverting Eq. (9). The slight difference between the parameters $\alpha$, $\tau$ and $n$ reconstructed from the rate and the original ones for $L_{\bar{\nu}_{e}}$ can be explained by the weak time dependence of $\langle E_{\nu}\rangle$, $\langle E_{\nu}^{2}\rangle$, and $\beta_{\nu}$ entering in the computation of the rate. Therefore, one can fit the detected event rate with the same functional form used for the luminosity, and from the reconstructed fitting parameters one can get information on the PNS mass and on the EoS, as discussed in the previous Sections. However, the accurate reconstruction of the parameters is beyond the scope of this analysis and will be the subject of future work dealing with their possible reconstruction using current and future neutrino detectors.
@@ VIII Conclusions
In this paper we have investigated whether the simple analytical function of Eq. (2) can be used as a parametric fit to the SN neutrino luminosities during the Kelvin-Helmholtz cooling phase of the PNS. For this purpose we have considered a set of several 1D simulations for different NS masses and EoSs. Our benchmark models account for PNS convection, which has a strong impact on the cooling evolution and its associated neutrino signal.
@@
Our fit function employs four free parameters, namely a normalization factor $C$, a power-law exponent $\alpha$ for the time, and exponential cooling time scale $\tau$, and an exponent $n$ of $(t/\tau)$ in the exponential function. Their characteristic dependence on the PNS mass and on the EoS can be used to draw inferences on these latter properties, if the parameters are deduced from the neutrino signal of a future Galactic SN explosion. For this purpose, we plan to investigate in future work how one can infer the parameter values of the neutrino luminosity from the SN neutrino signal measured in large underground detectors. We have demonstrated that this possibility is facilitated by the fact that the time evolution of the detected event rate depends on the neutrino luminosity $L_{\nu}$, the average neutrino energy $\langle E_{\nu}\rangle$, and the rms energy $\sqrt{\langle E_{\nu}^{2}\rangle}$, but the time dependence of $\langle E_{\nu}\rangle$ and $\sqrt{\langle E_{\nu}^{2}\rangle}$ is weak. This allowed us to show that Eq. (2) provides a good functional form to also fit the time evolution of the observed neutrino signal. Therefore, for a first estimation of the $L_{\nu}$-fit parameters from the event rate measured by a SN neutrino detector, one can simply apply our analytical expression for the luminosity fit and make use of the assumption that the average neutrino energy and the pinching parameter characterizing the spectral shape are constant in time.
@@
A number of caveats of our study reported here need to be mentioned. First, the present analysis and our proposed luminosity fit are based on 1D SN and PNS cooling simulations using a fairly limited set of cases for the NS EoS. The general applicability of the fit function of Eq. (2) needs confirmation by testing a much larger variety of EoS models with a wide range of fundamental nuclear physics inputs that are compatible with all experimental, theoretical, and astrophysical constraints on the properties of nuclear matter and observed NSs. In particular, possible correlations of some of the fitting parameters [e.g., the relation in Eq. (5)] require confirmation based on a wider spectrum of nuclear EoS representations. Second, our 1D SN and PNS calculations disregard 3D effects such as long-lasting accretion onto the PNS (continuing also after the onset of the explosion, when in 1D models accretion abruptly stops) and fallback of some initial explosion ejecta during the late PNS evolution [31, 16, 32]. Moreover, the mixing-length treatment of PNS convection in our 1D models will have to be validated by long-time 3D simulations of PNS cooling once such calculations with good spatial resolution become available. In particular, this will also provide a test whether the fit function of Eq. (2) and our best-fit parameter values are compatible with 3D results for PNS cooling. If so, any deviation from the luminosity evolution described by our fit function would signal additional contributions to the neutrino emission added on top of the cooling component from the PNS. Thus our $L_{\nu}$-fit could help to diagnose, disentangle, and describe such secondary neutrino emission phenomena in the neutrino measurement for a future Galactic SN. Finally, it will have to be seen how our fitting function reacts to additional, so far disregarded effects of potential importance in neutrino-cooling SN cores, for example fast flavor conversion of neutrino-antineutrino pairs, which could have a major impact on the neutrino emission properties [33, 34], or extra cooling associated with the emission of light, weakly interacting beyond-standard-model particles (e.g., axions [35]). Again, our neutrino luminosity fits could help to diagnose such effects beyond current standard SN modeling, once simulations including this new physics become available to be analyzed for long-time fitting.
@@
In conclusion, we are confident that the simplicity of our fitting procedure will make it a useful tool for the neutrino community to describe the SN neutrino signal expected in a high-statistics detection, to probe a future SN neutrino measurement, and to infer valuable information on the PNS mass, nuclear EoS, and different signal components.
@@ Appendix A Tables for the final-time parameters
Here we provide the values of the final simulation time $t_{\rm fin}$, the time $t_{\nu_{i},{\rm max}}$ when the quantity $tL_{\nu_{i}}$ reaches its maximum, $t_{\nu_{i},{\rm max}}\,L_{\nu_{i}}(t_{\nu_{i},{\rm max}})\equiv t_{\nu_{i},\rm max}L_{\nu_{i},\,\rm max}$, and the selected cut time $t_{\nu_{i},{\rm c}}$, for neutrinos in Tab. 2 and for antineutrinos in Tab. 3 for all the simulations considered in our study. For each (anti)neutrino species, the selected cut time $t_{\nu_{i},{\rm c}}$ is the time when the quantity $t\,L_{\nu_{i}}$ is reduced relative to its maximum value by a factor $X_{\nu_{i}}^{\rm c}=t_{\nu_{i},\rm c}L_{\nu_{i},\,\rm c}/t_{\nu_{i},\rm max}L_{\nu_{i},\,\rm max}=0.15$, in order to take a common final time for all neutrino species. As shown by the quantities in bold print in Tab. 2 and Tab. 3, $t_{\nu_{i},{\rm c}}>t_{\rm fin}$ for 1.62-SFHx-c and 1.61-LS220-c, i.e. these simulations stop before reaching $t_{\nu_{i},{\rm c}}$. Therefore, to check the impact of convection in these two cases we cut our simulation outputs at earlier times, $t_{\nu_{i},{\rm Co}}$, when $X_{\nu_{i}}^{\rm Co}=t_{\nu_{i},{\rm Co}}L_{\nu_{i},\,{\rm Co}}/t_{\nu_{i},\rm max}L_{\nu_{i},\,\rm max}=0.22$ for simulations with SFHx and $X_{\nu_{i}}^{{\rm Co}}=0.19$ for LS220. We show $t_{\rm fin}$, $t_{\nu_{i},{\rm max}}$ and $t_{\nu_{i},{\rm Co}}$ used to check the impact of convection for electron and muon (anti)neutrinos in Tab. 4.
@@ Appendix B Tables for best-fit parameters of luminosities for neutrinos and antineutrinos of all flavors
with $C$, $\alpha$, $\tau$ and $n$ being free parameters. We show values of the fit parameters for $\nu_{e}$ and $\bar{\nu}_{e}$ in Tab. 5, for $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ in Tab. 6 and for $\nu_{\tau}$ and $\bar{\nu}_{\tau}$ in Tab. 7, obtained with the NonlinearModelFit function in Mathematica.
@@ Appendix C Tables for the linear relations between luminosity-fitting parameters and PNS mass
In Tab. 8 we provide the best-fit values and their $1\,\sigma$ errors for the parameters $K_{0}$ and $K_{1}$ that describe the linear dependencies of the parameter values in the $L_{\nu}$-fit of Eq. (2) on the PNS mass $M_{\rm NS}$ at fixed EoS, for all neutrino and antineutrino species:
@@
where $K=C,\,\alpha,\,\tau,\,n$. The larger values of the relative uncertainties on $K_{0}$ and $K_{1}$ in Tab. 8 and the widths of the confidence bands in Figs. 8, 9, and 10 suggest that the linear fits work better for Class A EoSs than for LS220. At fixed EoS (in particular for Class A EoSs), the linear fits are excellent for $C$ and $\tau$ and slightly worse for $\alpha$ and $n$, featuring larger relative errors of the best-fit values of the parameters in Eq. (12) (see, e.g., the values of $\alpha_{1}$ and $n_{1}$).

### v3

@@ VII Counting rate in neutrino detectors
In SK, the main detection process is inverse $\beta$ decay, $\bar{\nu}_{e}\,p\to n\,e^{+}$, where the final-state positron shows up by its Cherenkov radiation. Because of the similarity of the electron and non-electron antineutrino luminosities and spectra in our models during PNS Kelvin-Helmholtz cooling at times $t\gtrsim 1$ s after bounce, flavor conversions are not a major effect and can be neglected in our simplified analysis. Therefore the expected rate can be written as
@@
where $N_{p}=1.51\times 10^{33}$ is the number of protons for a 22.5 kton Cherenkov detector. Here, we follow Ref. [58] for the limits of integration in $dE_{\nu}$ and we integrate the positron energies above the energy threshold $E_{\rm th,SK}=5~{}\operatorname{MeV}$. We mention that SK is essentially background free. Estimates for the future Hyper-Kamiokande detector with fiducial mass of 187 kton [59] can be obtained by rescaling the counting rate computed for SK by a factor $\sim 8.3$, without affecting the temporal evolution of the signal.
Figure 15: Left panel: Simulation results for the average energy $\langle E_{\bar{\nu}_{e}}\rangle$ (blue), rms energy $\sqrt{\langle E_{\bar{\nu}_{e}}^{2}\rangle}$ (orange) and spectral shape parameter $\beta_{\bar{\nu}_{e}}$ (green) from model 1.62-DD2. Right panel: $\bar{\nu}_{e}$-induced counting rate $tR_{{\bar{\nu}}_{e}}$ in SK computed from simulation data (blue), the fit of the counting rate (orange), and the fit of $tL_{{\bar{\nu}}_{e}}$ (dashed line).
To obtain a numerical estimate, we use data from model 1.62-DD2 as an example, but the same analysis is valid for all of the models. In the left panel of Fig. 15 we plot the average energy $\langle E_{\bar{\nu}_{e}}\rangle$, the root-mean-square (rms) energy $\sqrt{\langle E_{\bar{\nu}_{e}}^{2}\rangle}$, and the shape parameter $\beta_{\bar{\nu}_{e}}$ as a function of the time, between 1 s and $t_{\bar{\nu}_{e},{\rm c}}=7.94\,\operatorname{s}$. These quantities exhibit a weak time dependence, with $\langle E_{\bar{\nu}_{e}}\rangle\approx(12-13.5)\,\operatorname{MeV}$, $\sqrt{\langle E_{\bar{\nu}_{e}}^{2}\rangle}\approx(14-15)\,\operatorname{MeV}$ and $\beta_{\bar{\nu}_{e}}\approx 2.5-3$ at $t\lesssim 5\,\operatorname{s}$ and a decrease at later times. Here we focus on the 1.62-DD2 simulation as a representative case, but qualitatively similar results can be obtained with all the other models. Basic information on the time evolution of the mean neutrino energies for all considered EoSs can be found in Ref. [43] (see Fig. 3 therein, where the time evolution of the average neutrino energies for the $1.44~{}M_{\odot}$ with different nuclear EoSs are shown).
The rate in SK can be simply estimated as
$R_{\bar{\nu}_{e}}\approx\frac{L_{\bar{\nu}_{e}}}{4\pi\,D^{2}\langle E_{\bar{\nu}_{e}}\rangle}\,N_{\rm SK}\,\langle\sigma\rangle\,,$ (9)
@@
dependent on $\langle E_{\bar{\nu}_{e}}\rangle$ and $\beta_{\bar{\nu}_{e}}$ [and also on $\sqrt{\langle E_{\bar{\nu}_{e}}^{2}\rangle}$ via Eq. (7)]. Therefore, we expect that the observed rate $R_{\bar{\nu}_{e}}$ will follow the time dependence of the neutrino luminosity because of the weak time dependence of the other parameters, and that we can also fit it with our analytical formula in Eq. (2). In the right panel of Fig. 15 we show the quantity $tR_{\bar{\nu}_{e}}$ obtained from Eq. (8) using data from model 1.62-DD2 (blue line), its fit (orange), and the fitted $tL_{{\bar{\nu}}_{e}}$ (dashed line). As shown by the best-fit parameters for $t\,R_{\bar{\nu}_{e}}$ given in the plot, the fit leads to $C=8.28\times 10^{2}\,\operatorname{s}^{-1}$, $\alpha=0.54$, $\tau=6.17\,\operatorname{s}$ and $n=4.61$, to be compared with the ones for $tL_{\bar{\nu}_{e}}$. The parameters $\alpha_{\bar{\nu}_{e}}$, $\tau_{\bar{\nu}_{e}}$ and $n_{\bar{\nu}_{e}}$ are well reconstructed, while one could get information on the normalization factor $C_{\bar{\nu}_{e}}$ by inverting Eq. (9). The slight difference between the parameters $\alpha$, $\tau$ and $n$ reconstructed from the rate and the original ones for $L_{\bar{\nu}_{e}}$ can be explained by the weak time dependence of $\langle E_{\nu}\rangle$, $\langle E_{\nu}^{2}\rangle$, and $\beta_{\nu}$ entering in the computation of the rate. Therefore, one can fit the detected event rate with the same functional form used for the luminosity, and from the reconstructed fitting parameters one can get information on the PNS mass and on the EoS, as discussed in the previous Sections. However, the accurate reconstruction of the parameters is beyond the scope of this analysis and will be the subject of future work dealing with their possible reconstruction using current and future neutrino detectors. There, more comprehensive information on the time evolution of $\langle E_{\nu}\rangle$, $\langle E_{\nu}^{2}\rangle$ and $\beta_{\nu}$ will be provided.
@@ VIII Conclusions
In this paper we have investigated whether the simple analytical function of Eq. (2) can be used as a parametric fit to the SN neutrino luminosities during the Kelvin-Helmholtz cooling phase of the PNS. For this purpose we have considered a set of several 1D simulations for different NS masses and EoSs. Our benchmark models account for PNS convection, which has a strong impact on the cooling evolution and its associated neutrino signal. In particular, we presented fits for the time-dependent neutrino luminosities from numerical PNS cooling simulations and reported systematic dependences of the fitting-parameter values as functions of the NS mass. Future work is desirable where these fits are connected to analytic descriptions and basic PNS and EoS properties in a more formal way, e.g., similar to what was done for PNS cooling models that did not include the effects of PNS convection (see, e.g., [41, 60]).
@@
Our fit function employs four free parameters, namely a normalization factor $C$, a power-law exponent $\alpha$ for the time, an exponential cooling timescale $\tau$, and an exponent $n$ of $(t/\tau)$ in the exponential function. Their characteristic dependence on the PNS mass and on the EoS can be used to draw inferences on these latter properties, if the parameters are deduced from the neutrino signal of a future Galactic SN explosion. For this purpose, we plan to investigate in future work how one can infer the parameter values of the neutrino luminosity from the SN neutrino signal measured in large underground detectors. We have demonstrated that this possibility is facilitated by the fact that the time evolution of the detected event rate depends on the neutrino luminosity $L_{\nu}$, the average neutrino energy $\langle E_{\nu}\rangle$, and the rms energy $\sqrt{\langle E_{\nu}^{2}\rangle}$, but the time dependence of $\langle E_{\nu}\rangle$ and $\sqrt{\langle E_{\nu}^{2}\rangle}$ is weak. This allowed us to show that Eq. (2) provides a good functional form to also fit the time evolution of the observed neutrino signal. Therefore, for a first estimation of the $L_{\nu}$-fit parameters from the event rate measured by a SN neutrino detector, one can simply apply our analytical expression for the luminosity fit and make use of the assumption that the average neutrino energy and the pinching parameter characterizing the spectral shape are constant in time.
@@
A number of caveats of our study reported here need to be mentioned. First, the present analysis and our proposed luminosity fit are based on 1D SN and PNS cooling simulations using a fairly limited set of cases for the NS EoS. The general applicability of the fit function of Eq. (2) needs confirmation by testing a much larger variety of EoS models with a wide range of fundamental nuclear physics inputs that are compatible with all experimental, theoretical, and astrophysical constraints on the properties of nuclear matter and observed NSs. In particular, possible correlations of some of the fitting parameters [e.g., the relation in Eq. (5)] require confirmation based on a wider spectrum of nuclear EoS representations. Second, our 1D SN and PNS calculations disregard 3D effects such as long-lasting accretion onto the PNS (continuing also after the onset of the explosion, when in 1D models accretion abruptly stops) and fallback of some initial explosion ejecta during the late PNS evolution [61, 43, 62]. Moreover, the mixing-length treatment of PNS convection in our 1D models will have to be validated by long-time 3D simulations of PNS cooling once such calculations with good spatial resolution become available. In particular, this will also provide a test whether the fit function of Eq. (2) and our best-fit parameter values are compatible with 3D results for PNS cooling. If so, any deviation from the luminosity evolution described by our fit function would signal additional contributions to the neutrino emission added on top of the cooling component from the PNS. Thus, our $L_{\nu}$-fit could help to diagnose, disentangle, and describe such secondary neutrino emission phenomena in the neutrino measurement for a future Galactic SN. Finally, it will have to be seen how our fitting function reacts to additional, so far disregarded effects of potential importance in neutrino-cooling SN cores, for example fast flavor conversion of neutrino-antineutrino pairs, which could have a major impact on the neutrino emission properties [63, 64], or extra cooling associated with the emission of light, weakly interacting beyond-standard-model particles (e.g., axions [65]). Again, our neutrino luminosity fits could help to diagnose such effects beyond current standard SN modeling, once simulations including this new physics become available to be analyzed for long-time fitting.
@@
In conclusion, we are confident that the simplicity of our fitting procedure will make it a useful tool for the neutrino community to describe the SN neutrino signal expected in a high-statistics detection, to probe a future SN neutrino measurement, and to infer valuable information on the PNS mass, nuclear EoS, and different signal components (see Ref. [66] for a recent approach in this direction).
The considered model results are adopted from Ref. [43] and are available in the Garching Core-collapse Supernova Archive [49] upon request.
@@ Appendix A Tables for the final-time parameters
Here we provide the values of the final simulation time $t_{\rm fin}$, the time $t_{\nu_{i},{\rm max}}$ when the quantity $tL_{\nu_{i}}$ reaches its maximum, $t_{\nu_{i},{\rm max}}\,L_{\nu_{i}}(t_{\nu_{i},{\rm max}})\equiv t_{\nu_{i},\rm max}L_{\nu_{i},\,\rm max}$, and the selected cut time $t_{\nu_{i},{\rm c}}$, for neutrinos in Table 2 and for antineutrinos in Table 3 for all the simulations considered in our study. For each (anti)neutrino species, the selected cut time $t_{\nu_{i},{\rm c}}$ is the time when the quantity $t\,L_{\nu_{i}}$ is reduced relative to its maximum value by a factor $X_{\nu_{i}}^{\rm c}=t_{\nu_{i},\rm c}L_{\nu_{i},\,\rm c}/t_{\nu_{i},\rm max}L_{\nu_{i},\,\rm max}=0.15$, in order to take a common final time for all neutrino species. As shown by the quantities in bold print in Table 2 and Table 3, $t_{\nu_{i},{\rm c}}>t_{\rm fin}$ for 1.62-SFHx-c and 1.61-LS220-c, i.e. these simulations stop before reaching $t_{\nu_{i},{\rm c}}$. Therefore, to check the impact of convection in these two cases we cut our simulation outputs at earlier times, $t_{\nu_{i},{\rm Co}}$, when $X_{\nu_{i}}^{\rm Co}=t_{\nu_{i},{\rm Co}}L_{\nu_{i},\,{\rm Co}}/t_{\nu_{i},\rm max}L_{\nu_{i},\,\rm max}=0.22$ for simulations with SFHx and $X_{\nu_{i}}^{{\rm Co}}=0.19$ for LS220. We show $t_{\rm fin}$, $t_{\nu_{i},{\rm max}}$ and $t_{\nu_{i},{\rm Co}}$ used to check the impact of convection for electron and muon (anti)neutrinos in Table 4.
@@ Appendix B Tables for best-fit parameters of luminosities for neutrinos and antineutrinos of all flavors
with $C$, $\alpha$, $\tau$ and $n$ being free parameters. We show values of the fit parameters for $\nu_{e}$ and $\bar{\nu}_{e}$ in Table 5, for $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ in Table 6 and for $\nu_{\tau}$ and $\bar{\nu}_{\tau}$ in Table 7, obtained with the NonlinearModelFit function in Mathematica.
@@ Appendix C Tables for the linear relations between luminosity-fitting parameters and PNS mass
In Table 8 we provide the best-fit values and their $1\,\sigma$ errors for the parameters $K_{0}$ and $K_{1}$ that describe the linear dependencies of the parameter values in the $L_{\nu}$-fit of Eq. (2) on the PNS mass $M_{\rm NS}$ at fixed EoS, for all neutrino and antineutrino species:
@@
where $K=C,\,\alpha,\,\tau,\,n$. The larger values of the relative uncertainties on $K_{0}$ and $K_{1}$ in Table 8 and the widths of the confidence bands in Figs. 8, 9, and 10 suggest that the linear fits work better for Class A EoSs than for LS220. At fixed EoS (in particular for Class A EoSs), the linear fits are excellent for $C$ and $\tau$ and slightly worse for $\alpha$ and $n$, featuring larger relative errors of the best-fit values of the parameters in Eq. (12) (see, e.g., the values of $\alpha_{1}$ and $n_{1}$).